# BugMagnifier: TON Transaction Simulator for Revealing Smart Contract Vulnerabilities

### v1

@@ Abstract
The Open Network (TON) blockchain employs an asynchronous execution model that introduces unique security challenges for smart contracts, particularly race conditions arising from unpredictable message processing order. While previous work established vulnerability patterns through static analysis of audit reports, dynamic detection of temporal dependencies through systematic testing remains an open problem. We present BugMagnifier, a transaction simulation framework that systematically reveals vulnerabilities in TON smart contracts through controlled message orchestration. Built atop TON Sandbox and integrated with the TON Virtual Machine (TVM), our tool combines precise message queue manipulation with differential state analysis and probabilistic permutation testing to detect asynchronous execution flaws. Experimental evaluation demonstrates BugMagnifier’s effectiveness through extensive parametric studies on purpose-built vulnerable contracts, revealing message ratio-dependent detection complexity that aligns with theoretical predictions. This quantitative model enables predictive vulnerability assessment while shifting discovery from manual expert analysis to automated evidence generation. By providing reproducible test scenarios for temporal vulnerabilities, BugMagnifier addresses a critical gap in the TON security tooling, offering practical support for safer smart contract development in asynchronous blockchain environments.
@@ I Introduction
Blockchain architectures face an inherent tension between transactional atomicity and network scalability, a fundamental trade-off defining platform security profiles [1, 2, 3]. Ethereum’s synchronous model [4] demonstrates this through complex atomic operations such as decentralized exchange (DEX) arbitrage [5, 6], enabling coordinated asset swaps within single transactions. These operations rely on deterministic execution ordering, ensuring static market conditions during processing. However, this atomicity restricts throughput to approximately 30 transactions per second (TPS), limiting scalability to single-chain performance.
The Open Network (TON) addresses scalability through a sharded architecture where transactions decompose into independent messages processed concurrently across chains [7, 8]. This asynchronous model enables theoretical throughput exceeding 100,000 TPS, i.e., an improvement of over three orders of magnitude. However, this innovation introduces novel security challenges. Where Ethereum implements arbitrage atomically, TON requires decomposition into interdependent messages. Developers must manage state changes between executions, cross-shard fees, and partial failure recovery. Unpredictable message interleaving creates distinct vulnerability patterns, with temporal dependencies becoming predominant.
TON’s architecture inherently amplifies race condition risks through five-phase transaction processing , where intermediate states become visible to subsequent messages. The cell-based storage of the TVM exacerbates risks through dynamic fee calculations based on the depth of the cell tree, enabling financial attacks via strategic message sequences [9]. Official TON security guidelines explicitly warn about race conditions in message flows and account destruction scenarios, categorizing them as critical vulnerabilities requiring preventive measures.
While Ethereum’s ecosystem has experienced critical vulnerabilities like The DAO’s reentrancy attack and Parity’s access control flaws [10, 11, 12], these differ fundamentally from TON’s architecturally induced race conditions. Ethereum’s documented race condition vulnerabilities [13] typically require deliberate transaction ordering rather than emerging from inherent design constraints. TON’s immature security ecosystem compounds these challenges; existing analyzers like TONScanner [14] and TSA [15] focus on code defects rather than temporal vulnerabilities, leaving developers without systematic methods to detect message-ordering flaws.
@@
    Designed and developed the TON-Complete Transaction Simulation that integrates with TON Sandbox to enable local emulation of TON’s complete transaction processing cycle, including all five execution phases, a capability absent in existing tools that typically support only the compute phase.
@@ II-A TON Virtual Machine (TVM)
The TON Virtual Machine (TVM) serves as the deterministic execution environment for smart contracts on The Open Network, operating as a stack-based virtual machine with specific architectural features tailored to TON’s sharded blockchain design [16]. As blockchains operate in an untrusted environment where participants must self-verify computational consistency, computation results need to be hardware-independent; virtual machines like TVM are the standard choice to ensure this determinism across all network nodes. Unlike Ethereum’s virtual machine (EVM), which processes transactions as atomic units, TVM is specifically engineered to handle TON’s message-centric execution model, where transactions decompose into independent messages processed across (parallel) shard chains. The TVM state comprises five critical components: Stack (for data manipulation), Control Registers (managing execution flow), Current Continuation (tracking program state), Current Codepage (storing executable code), and Gas Limits (enforcing computational constraints).
@@
The TVM instruction set incorporates blockchain-specific functionality through application-oriented primitives, including message handling (SEND_RAW_MESSAGE), gas management (ACCEPT_MSG_VALUE), and cryptographic operations (SHA256_C). Notably, TVM implements a five-phase transaction processing model: storage, credit, compute, action, and bounce, each with distinct gas accounting rules and state transition semantics. This multi-phase execution model creates unique security challenges, as vulnerabilities can manifest across phase boundaries rather than within isolated transaction contexts. The compute phase, where contract logic is executed, operates under strict gas limits that vary based on preceding phase outcomes, requiring precise gas calculation to avoid out-of-gas failures that could leave contracts in inconsistent states.
@@ II-B Asynchronous Transaction Execution
TON’s asynchronous execution model fundamentally differs from conventional blockchains through its message-driven architecture, where transactions decompose into independent messages processed concurrently across shard chains. Each message undergoes five sequential processing phases: storage (calculating storage fees), credit (accounting for incoming funds), compute (executing contract logic), action (generating new messages), and bounce (handling failed message processing). This phased execution enables higher throughput by allowing parallel processing of independent messages while maintaining deterministic outcomes through the consensus algorithm that establishes a final total order for messages.
The asynchronous model introduces critical security considerations absent in synchronous blockchains. While the consensus algorithm ensures a final deterministic outcome across the network, the concurrent processing of messages across shards creates a period of temporal non-determinism from the perspective of a single smart contract. The order in which messages from different senders interleave in their queue is not fixed until included in a block, creating race conditions when multiple messages target the same contract state. Unlike Ethereum’s transaction atomicity, where state changes occur only upon successful transaction completion, TON contracts experience intermediate state transitions after each message processing phase, particularly during the compute phase, which can be observed by subsequent messages. This behavior enables temporal vulnerabilities where the contract state appears consistent during individual message processing but becomes inconsistent when messages interleave unexpectedly.
@@ III Related Work
Unlike the well-studied Ethereum ecosystem, smart contract security in TON remains an under-researched area, primarily due to the platform’s architectural divergence from conventional blockchain models. TON’s asynchronous execution model, sharded architecture, and cell-based data structures introduce unique security challenges that render many Ethereum-focused analysis techniques ineffective. The fundamental difference lies in TON’s message-driven execution paradigm, where transactions decompose into independent messages processed across parallel chains, creating vulnerability patterns fundamentally distinct from those observed in synchronous blockchains.
To address this gap, Song et al. [14] conducted a pioneering study identifying eight key vulnerability types specific to TON, Bad Randomness, Precision Loss, Unchecked Return, Global Var Redefined, Improper Function Modifier, Unchecked Bounced Message, Inconsistent Data, Lack End Parse, and proposed TONScanner, a static analysis framework designed to detect these defects. TONScanner leverages the FunC compiler’s frontend to transform FunC source code into an intermediate representation (IR) structured as a directed acyclic graph (DAG). The tool then constructs a control flow graph (CFG) and converts it into static single assignment (SSA) form to streamline analysis. To accommodate TON’s unique data structures, TONScanner integrates specialized components such as Data Dependency Analysis, Call Graph Construction, Taint Analysis, and Cell Construct Analysis. These modules collectively enable the detection of the eight identified defect categories.
Technically, TONScanner’s Cell Construct Analysis module addresses TON’s distinctive cell-based storage mechanism by implementing a recursive cell tree validator that checks for proper cell serialization and deserialization patterns. This component specifically targets vulnerabilities arising from improper cell handling, such as cell overflow conditions and incorrect bit string manipulations, issues that are nonexistent in Ethereum’s account-based model. The tool’s Taint Analysis engine has been extended with TON-specific sinks, including message sending operations (SEND_RAW_MESSAGE) and bounce handling mechanisms (SET_BOUNCE), which are critical for identifying message redirection vulnerabilities.
In an evaluation of 1,640 real-world TON smart contracts, TONScanner uncovered 14,995 defects, highlighting the prevalence of developer errors in the ecosystem. A manual audit of randomly sampled results confirmed the tool’s 97.49% precision, demonstrating its reliability for practical use. Notably, the study revealed that Unchecked Bounced Message vulnerabilities accounted for 38.7% of all detected issues, underscoring the significance of TON’s unique bounce phase in security analysis, a vulnerability class absent in Ethereum ecosystems.
A more recent advancement in TON smart contract analysis is the Ton Symbolic Analyzer (TSA) [15], which operates at the bytecode level using symbolic virtual machines and satisfiability modulo theories (SMT) solvers. This tool is specifically designed to: (1) detect potential runtime errors in the TON Virtual Machine (TVM), (2) generate regression tests based on Blueprint framework specifications, and (3) identify malicious honeypot contracts, a class of deceptive smart contracts engineered to exploit users. TSA demonstrates notable effectiveness in flagging fraudulent contract behavior, verifying expected program execution paths, and detecting anomalous runtime conditions.
TSA’s technical innovation lies in its TVM-aware symbolic execution engine, which models TON’s five-phase transaction processing (storage, credit, compute, action, and bounce) as distinct symbolic states. This approach enables the tool to detect vulnerabilities that span multiple transaction phases, particularly critical for identifying race conditions that manifest across message boundaries. The implementation incorporates a custom SMT theory for TON’s cell operations, allowing precise modeling of cell serialization vulnerabilities that would be undetectable through conventional symbolic execution. The evaluation [15] demonstrated TSA’s capability to identify complex multi-phase vulnerabilities with 89.3% recall, though at significant computational cost (average 27 minutes per contract).
While existing tools represent important initial efforts in addressing TON-specific vulnerabilities, they lack the comprehensive coverage and maturity of their Ethereum counterparts. As demonstrated in our previous work [17], which established a detailed taxonomy of TON smart contract vulnerabilities and proposed a structured audit methodology, there remains a significant gap in practical analysis tools for this emerging ecosystem. Most critically, current solutions fail to address the temporal dimension of TON security, specifically, the ability to systematically explore message ordering permutations and their impact on contract state. Static analysis tools such as TONScanner cannot model runtime message interleaving, while symbolic execution approaches like TSA become computationally intractable when analyzing multiple message sequences.
In the present study, we introduce BugMagnifier, an advanced debugging instrument designed specifically for TON smart contract analysis. This tool complements our earlier theoretical framework by providing: (1) dynamic analysis capabilities through precise message queue manipulation, (2) enhanced vulnerability detection mechanisms for temporal vulnerabilities via message order permutations, and (3) real-time debugging features tailored to TON’s unique architecture, including full transaction phase visibility and cell-level state inspection. Unlike existing solutions, BugMagnifier operates at the message granularity level, enabling systematic exploration of execution paths that span multiple transactions, a capability essential for identifying the temporal vulnerabilities that dominate TON’s security landscape.
Through this combined theoretical and practical approach, we aim to bridge the gap between Ethereum’s mature audit methodologies and the evolving needs of the TON ecosystem. BugMagnifier specifically addresses the critical deficiency in current tooling by providing the first practical solution for systematic race condition detection in TON smart contracts, a vulnerability class responsible for the majority of critical security incidents in the ecosystem.

### v2

@@ Abstract
The Open Network (TON) blockchain employs an asynchronous execution model that introduces unique security challenges for smart contracts. A primary concern is race conditions arising from unpredictable message processing order. While previous work established vulnerability patterns through static analysis of audit reports, dynamic detection of temporal dependencies through systematic testing remains an open problem. This study proposes a dynamic evaluation methodology based on controlled message orchestration to systematically expose vulnerabilities in asynchronous smart contracts. By synthesizing precise message queue manipulation with differential state analysis and probabilistic permutation testing, we establish a framework (viz, BugMagnifier) for identifying execution flaws that static methods miss. Experimental evaluation demonstrates BugMagnifier’s effectiveness through extensive parametric studies on purpose-built vulnerable contracts, revealing message ratio-dependent detection complexity that aligns with theoretical predictions. This quantitative model enables predictive vulnerability assessment while shifting discovery from manual expert analysis to automated evidence generation. By providing reproducible test scenarios for temporal vulnerabilities, BugMagnifier addresses a critical gap in the TON security tooling, offering practical support for safer smart contract development in asynchronous blockchain environments.
@@ I Introduction
Blockchain architectures face an inherent tension between transactional atomicity and network scalability, a fundamental trade-off defining platform security profiles [Buterin2021a, Kruglik2019a, Amelin2021]. Ethereum’s synchronous model [Buterin2014] demonstrates this through complex atomic operations such as decentralized exchange (DEX) arbitrage [Gramlich2024, Vostrikov2025], enabling coordinated asset swaps within single transactions. These operations rely on deterministic execution ordering, ensuring static market conditions during processing. However, this atomicity restricts throughput to approximately 30 transactions per second (TPS), limiting scalability to single-chain performance.
The Open Network (TON) addresses scalability through a sharded architecture where transactions decompose into independent messages processed concurrently across chains [Durov2019, Berger2019]. This asynchronous model enables theoretical throughput exceeding 100,000 TPS, i.e. an improvement of more than three orders of magnitude. However, this innovation introduces novel security challenges. Where Ethereum implements arbitrage atomically, TON requires decomposition into interdependent messages. Developers must manage state changes between executions, cross-shard fees, and partial failure recovery. Unpredictable message interleaving creates distinct vulnerability patterns, with temporal dependencies becoming predominant.
TON’s architecture inherently amplifies race condition risks through five-phase transaction processing, where intermediate states become visible to subsequent messages. Cell-based storage of TVM exacerbates risks through dynamic fee calculations based on the depth of the cell tree, enabling financial attacks via strategic message sequences [TON2025a]. Official TON security guidelines explicitly warn about race conditions in message flows and account destruction scenarios, categorizing them as critical vulnerabilities requiring preventive measures.
Whilst Ethereum’s ecosystem has experienced critical vulnerabilities like The DAO’s reentrancy attack and Parity’s access control flaws [Atzei2017, Chen2022a, Soud2024], these differ fundamentally from TON’s architecturally induced race conditions. Ethereum’s documented race condition vulnerabilities [Breidenbach2018] typically require deliberate transaction ordering rather than emerging from inherent design constraints. The nascent TON security ecosystem currently offers analyzers like TONScanner [Song2025] and TSA [Espirito2025], which are effective for identifying static code defects (e.g., bad randomness, unchecked returns). However, these tools operate via static analysis and symbolic execution, methodologies that are inherently unable to simulate the temporal execution paths arising from non-deterministic message interleaving. Consequently, a critical gap exists for detecting dynamic, state-dependent race conditions, leaving developers without systematic methods to validate contract behavior under asynchronous execution.
@@
    Designed and developed the TON-Complete Transaction Simulation that integrates with TON Sandbox to enable local emulation of the TON’s complete transaction processing cycle, including all five execution phases, a capability absent in existing tools that typically support only the compute phase.
@@ II-A TON Virtual Machine (TVM)
The TON Virtual Machine (TVM) serves as the deterministic execution environment for smart contracts on The Open Network, operating as a stack-based virtual machine with specific architectural features tailored to TON’s sharded blockchain design [Durov2020]. As blockchains operate in an untrusted environment where participants must self-verify computational consistency, computation results need to be hardware-independent; virtual machines like TVM are the standard choice to ensure this determinism across all network nodes. Unlike Ethereum’s virtual machine (EVM), which processes transactions as atomic units, TVM is specifically engineered to handle TON’s message-centric execution model, where transactions decompose into independent messages processed across (parallel) shard chains. The TVM state comprises five critical components: Stack (for data manipulation), Control Registers (managing execution flow), Current Continuation (tracking program state), Current Codepage (storing executable code), and Gas Limits (enforcing computational constraints).
@@
The TVM instruction set incorporates blockchain-specific functionality through application-oriented primitives, including message handling (SEND_RAW_MESSAGE), gas management (ACCEPT_MSG_VALUE), and cryptographic operations (SHA256_C). Notably, TVM implements a five-phase transaction processing model: storage, credit, compute, action, and bounce, each with distinct gas accounting rules and state transition semantics. This multi-phase execution model creates unique security challenges, as vulnerabilities can manifest across phase boundaries rather than within isolated transaction contexts. The compute phase, where contract logic is executed, operates under strict gas limits that vary based on preceding phase outcomes, requiring precise gas calculations to avoid out-of-gas failures that could leave contracts in inconsistent states.
@@ II-B Asynchronous Transaction Execution
TON’s asynchronous execution model fundamentally differs from conventional blockchains through its message-driven architecture, where transactions decompose into independent messages processed concurrently across shard chains. Each message undergoes five sequential processing phases: storage (calculating storage fees), credit (accounting for incoming funds), compute (executing contract logic), action (generating new messages) and bounce (handling failed message processing). This phased execution enables higher throughput by allowing parallel processing of independent messages while maintaining deterministic outcomes through the consensus algorithm that establishes a final total order for messages.
The asynchronous model introduces critical security considerations that are absent in synchronous blockchains. While the consensus algorithm ensures a final deterministic outcome across the network, the concurrent processing of messages across shards creates a period of temporal non-determinism from the perspective of a single smart contract. The order in which messages from different senders interleave in their queue is not fixed until included in a block, creating race conditions when multiple messages target the same contract state. Unlike Ethereum’s transaction atomicity, where state changes occur only upon successful transaction completion, TON contracts experience intermediate state transitions after each message processing phase, particularly during the compute phase, which can be observed by subsequent messages. This behavior enables temporal vulnerabilities where the contract state appears consistent during individual message processing but becomes inconsistent when messages interleave unexpectedly.
@@
A significant emerging trend is the application of Artificial Intelligence and Large Language Models (LLMs) for automated vulnerability discovery and exploit generation. Recent studies demonstrate that models like GPT-4 and Claude can analyze Solidity code and generate functional proof-of-concept exploits for known vulnerability patterns with notable success rates [Xiao2025]. Although these approaches excel at reasoning about code semantics and predicting flaws based on learned patterns, they operate through static analysis or limited symbolic execution. BugMagnifier represents a complementary paradigm: instead of *predicting* vulnerabilities from code, it *concretely executes* contract logic under systematically varied message orderings to *discover* temporal flaws. This dynamic exploration-based approach is crucial for detecting asynchronous race conditions that are often poorly represented in the training data of current LLMs.
@@ III Related Work
The security of smart contracts has matured into a significant research domain, though its focus remains predominantly on the Ethereum Virtual Machine (EVM) and its synchronous execution model. This section reviews the foundational work in smart contract vulnerability analysis, the tooling built for the EVM-centric paradigm, and the emerging research on transaction ordering in adversarial environments. We then contextualize existing TON-specific analysis within this broader landscape, concluding with a comparative analysis that highlights the gap filled by our work: a dynamic transaction simulator for the asynchronous TON blockchain.
III-A General Smart Contract Vulnerabilities and Empirical Studies
The systematic study of smart contract vulnerabilities was pioneered by researchers analyzing the EVM and Solidity. Early work established taxonomies of common flaws, such as reentrancy and integer overflows, which have become the primary targets for automated audit tools. This foundation is crucial, yet the specific vulnerability patterns are often tied to the semantics of the EVM opcodes and Solidity’s execution model.
Recent empirical studies continue to deepen this understanding by analyzing large-scale contract deployments. For instance, Chen et al. [Chen2019b] conducted a large-scale empirical study focusing on the critical task of identifying control flow in smart contracts, a foundational step for many advanced static analyzes. Their work on Ethereum contracts highlights the methodologies and challenges of reasoning about contract execution paths, which differs significantly in TON’s message-passing architecture. Furthermore, Li et al. [Li2025a] provided a contemporary empirical analysis of inconsistent state update vulnerabilities. While their study is grounded in the EVM paradigm, the core concept–unexpected state interleaving due to concurrent operations–is directly analogous to the race conditions endemic to TON’s asynchronous model, underscoring the universal nature of state consistency challenges in decentralized systems.
III-B Analysis Tools for the EVM and Synchronous Paradigms
The maturity of Ethereum security research is evidenced by a robust ecosystem of analysis tools employing static, dynamic, and formal methods. Static analysis frameworks like Slither [Feist2019] have set a high standard; by converting Solidity source code into an intermediate representation and applying data-flow and taint analysis, they efficiently detect dozens of common vulnerability patterns without execution. Automated invariant generation techniques [Liu2025a] further enhance this by inferring and checking the properties that must hold across all executions, providing a higher level of assurance for synchronous contract models. Similarity-based analysis [Chen2025] offers another angle by comparing contract codebases to known vulnerable patterns, efficiently scaling to large ecosystems. Dynamic analysis and fuzzing tools, such as Echidna [Grieco2020] and xFuzz [Xue2024], take a complementary approach by executing contracts with generated or mutated inputs to uncover unexpected runtime behavior. Dynamic exploit generation [Wang2022c] advances this further by automatically creating concrete proofs-of-concept for detected vulnerabilities. On the formal verification front, tools like the Certora Prover represent the high-assurance end of the spectrum, using mathematical proofs to guarantee that a contract’s implementation satisfies its formal specification. Similarly, correct-by-design approaches such as VeriSolid [Nelaturu2023] enable the specification and verification of interacting contracts’ state machines, offering strong guaranties for synchronous systems like Ethereum.
Whilst powerful, these tools are architected around the assumptions of the EVM: atomic transaction execution, a globally ordered block sequence, and synchronous state transitions. They reason about a single transaction’s effect on a contract’s state in isolation. Consequently, they lack the fundamental primitives to model TON’s core challenges: the non-deterministic interleaving of multiple independent messages targeting the same contract and the state visibility across the five distinct processing phases (storage, credit, compute, action, bounce). A tool like Slither cannot analyze the flow of messages between accounts, and a fuzzer like Echidna cannot systematically permute the order of a message queue to discover race conditions.
TABLE I: Comparative Analysis of Smart Contract Security Approaches
Tool / Approach Target Platform Core Technique Models Async. Queues? Primary Strength Limitation for TON Asynchronous Vulnerabilities
Slither [Feist2019] EVM Static Analysis × Fast detection of code pattern vulnerabilities in Solidity. EVM-centric; cannot model TON’s message lifecycle or cell semantics.
Echidna [Grieco2020] EVM Fuzzing × Property-based testing via execution with random inputs. Generates linear transaction sequences, not interleaved, concurrent messages.
MEV Research [Daian2020] EVM Economic Analysis * Models profit from adversarial transaction ordering. Focuses on miner economics, not contract-state race condition detection.
VeriSolid [Nelaturu2023] EVM Formal Verification / Model Checking × Correct-by-design construction and verification of interacting state machines. Designed for synchronous atomic transactions; cannot model TON’s message lifecycle or interleavings.
Concurrency Testing [Elmas2013, Terragni2025] Generic Software Dynamic Analysis (concepts) Systematic exploration of thread/interleaving bugs. Framework not adapted for blockchain message-passing or TVM state.
TONScanner [Song2025] TON Static Analysis × Detects 8+ TON-specific code defect patterns in FunC. No execution; cannot simulate message queues or detect ordering flaws.
TSA [Espirito2025] TON Symbolic Execution × Finds runtime errors Path explosion limits exploration of multi-message permutations.
BugMagnifier (Proposed) TON Dynamic Simulation & Permutation Testing Systematic race condition detection via controlled message orchestration. N/A (Designed to address the above limitations)
* Analyzes ordering within a synchronous block, not an asynchronous message queue.
III-C Adversarial Transaction Ordering and MEV Research
Research into Maximum Extractable Value (MEV) has extensively explored the adversarial manipulation of transaction ordering within a block, a concern that shares a conceptual lineage with TON’s race conditions. The seminal work by Daian et al. [Daian2020] in “Flash Boys 2.0” documented how bots can front-run, back-run, or sandwich transactions in Ethereum’s mempool to extract profit, leading to consensus instability and economic centralization. This body of work formally treats blockchain state as a function of an ordered transaction list and examines the consequences of perturbing that order.
Our work is philosophically aligned with this research in its focus on ordering-dependent state outcomes. However, a critical distinction exists: MEV research primarily studies the economic exploitation of ordering in a public mempool for profit within a synchronous block. In contrast, TON’s vulnerability is architectural: the lack of atomicity means that even two honest, correctly formatted messages can create a financially damaging race condition based solely on their arrival order, regardless of miner/validator intent. BugMagnifier applies the core investigative principle–systematically exploring ordering permutations–to this different, platform-inherent security problem.
III-D Foundations in Concurrent System Verification
The core challenge of detecting state-dependent race conditions in TON is conceptually analogous to the classic problem of verifying correctness in concurrent software, where non-deterministic thread scheduling can lead to Heisenbugs. Research in this domain has developed robust methodologies for systematic exploration. Frameworks like CONCURRIT introduced programmable thread interleaving for testing, allowing developers to specify and exhaustively explore specific scheduling scenarios of interest [Elmas2013]. Similarly, research on differential testing of concurrent classes provides formal methods for identifying behavioral divergences between different implementations under all possible interleavings [Terragni2025].
BugMagnifier adapts this foundational principle to the blockchain domain. It replaces the abstraction of thread scheduling with the concrete, controlled manipulation of TON’s message queue. While traditional tools reason about shared memory and locks, BugMagnifier operates on TVM state transitions and cell semantics. This connection positions our work not as an ad hoc TON-specific solution, but as a principled application of established concurrent system verification techniques to the novel execution model of asynchronous blockchains.
III-E Prior Work on TON Smart Contract Security
Security analysis for TON is a nascent field. Initial efforts have rightly focused on porting established techniques to the new platform. Song et al.’s TONScanner [Song2025] represents a critical first step, applying static analysis patterns to FunC source code to identify coding errors like unchecked bounced messages or precision loss. Similarly, the Ton Symbolic Analyzer (TSA) [Espirito2025] employs symbolic execution on TVM bytecode to find runtime errors. These tools are invaluable for detecting intra-transaction and single-message bugs, analogous to the first generation of Ethereum tools.
For Ethereum, formal verification tools like Certora Prover represent the high-assurance end of the security spectrum, using mathematical proofs to guarantee that a contract’s implementation satisfies its formal specification [Certora2025]. While such tools provide the strongest guaranties for synchronous blockchains, their adaptation to TON’s asynchronous model is an open challenge. Verifying temporal properties–such as the absence of race conditions across all possible message interleavings–requires sophisticated modeling of the message queue and TVM phases, a problem analogous to the formal verification of distributed, concurrent systems. BugMagnifier’s permutation testing can be viewed as a dynamic and practical stepping stone towards such formal guaranties, helping developers identify flawed contracts and gather concrete execution traces that could later inform the creation of formal models and specifications.
Their limitation, as highlighted in our introduction, is their inability to address the inter-message and temporal vulnerabilities that define TON’s unique threat model. Static analysis cannot simulate the execution of a message queue, and symbolic execution becomes computationally intractable when modeling all possible permutations of even a small set of concurrent messages. This creates a gap where a contract can pass all existing TON linters and verifiers, yet remain highly vulnerable to race conditions.
III-F Comparative Analysis and Identified Gap
As illustrated in Table I, existing tools are bifurcated into two categories: 1) mature solutions for a synchronous paradigm that cannot model TON architecture, and 2) nascent TON-specific tools that cannot analyze the temporal dependencies arising from concurrency. Research on transaction ordering, while insightful, does not provide a practical testing framework for TON smart contracts.
BugMagnifier is designed to bridge this gap directly. It takes the adversarial ordering exploration principle from MEV research and implements it as a practical, dynamic analysis tool for the TON blockchain. By providing granular control over the TVM’s execution phases and the message queue, it enables the systematic discovery of vulnerabilities that are invisible to all other approaches in the TON security toolkit. Our work thus advances the field by adapting a core security analysis methodology–permutation testing–to the unique challenges of asynchronous, sharded blockchain platforms.